OpenStudy (anonymous):

A stick 24 cm long is cut into two pieces such that one is 6 cm longer than the other. Find the length of the shorter piece.

OpenStudy (anonymous):

so Let length of one part Be X. then the other part will be X+6. (6 cm longer) and sum of both lengths is 24. so X + X +6 = 24; hence find the value of X

OpenStudy (anonymous):

24-6=X+X ?

OpenStudy (anonymous):

18= 2x and devide by x, x= 9

OpenStudy (anonymous):

24-9= 15 your two sizes are 15 and 9 cm